# LW-FedSSL: Resource-efficient Layer-wise Federated Self-supervised Learning

### v4

@@ Abstract
Many studies integrate federated learning (FL) with self-supervised learning (SSL) to take advantage of raw data distributed across edge devices. However, edge devices often struggle with high computational and communication costs imposed by SSL and FL algorithms. With the deployment of more complex and large-scale models, such as Transformers, these challenges are exacerbated. To tackle this, we propose the Layer-Wise Federated Self-Supervised Learning (LW-FedSSL) approach, which allows edge devices to incrementally train a small part of the model at a time. Specifically, in LW-FedSSL, training is decomposed into multiple stages, with each stage responsible for only a specific layer (or a block of layers) of the model. Since only a portion of the model is active for training at any given time, LW-FedSSL significantly reduces computational requirements. Additionally, only the active model portion needs to be exchanged between the FL server and clients, reducing the communication overhead. This enables LW-FedSSL to jointly address both computational and communication challenges in FL. Depending on the SSL algorithm used, it can achieve up to a 3.34×3.34$\times$3.34 × reduction in memory usage, 4.20×4.20$\times$4.20 × fewer computational operations (GFLOPs), and a 5.07×5.07$\times$5.07 × lower communication cost while maintaining performance comparable to its end-to-end training counterpart. Furthermore, we explore a progressive training strategy called Prog-FedSSL, which offers a 1.84×1.84$\times$1.84 × reduction in GFLOPs and a 1.67×1.67$\times$1.67 × reduction in communication costs while maintaining the same memory requirements as end-to-end training. While the resource efficiency of Prog-FedSSL is lower than that of LW-FedSSL, its performance improvements make it a viable candidate for FL environments with more lenient resource constraints.
@@ 1 Introduction
A significant portion of real-world data, valuable for practical AI applications, resides on edge devices. Such data is often unlabeled, making the adoption of self-supervised learning (SSL) strategies [1, 2, 3] essential. SSL enables models to learn from raw data by generating their own supervisory signals, minimizing the need for labeled datasets. Most SSL methods follow centralized training schemes, often requiring an exhaustive data collection process to build a centrally stored dataset. While some types of data, such as natural images, can be collected with relative ease, privacy-sensitive data—such as medical records—raise significant privacy concerns. Meanwhile, decentralized learning approaches like federated learning (FL) [4, 5] enable privacy-preserving collaborative model training. An FL system operates by collecting trained model parameters from edge devices (a.k.a., clients) instead of their data, thereby preserving data privacy. This mechanism helps safeguard the confidentiality of sensitive information held by different parties. As a result, federated self-supervised learning (FedSSL) has recently emerged as a promising approach for leveraging raw data in distributed environments [6, 7, 8].
Self-supervised learning (SSL), particularly with state-of-the-art models like Transformers [9, 10], imposes substantial computational demands. Unfortunately, clients in an FL environment often operate with limited computational and communication resources. Many clients may lack the necessary resources to participate in a conventional end-to-end FedSSL process. As a result, data from these clients is excluded from the training process, potentially degrading overall model performance. Therefore, it is crucial to explore resource-efficient FedSSL approaches that enable every FL client to participate in collaborative training. A straightforward way to reduce computational demands on low-memory devices is to conduct the training with a smaller batch size. However, using a small batch size in SSL can diminish the quality of learned representations and compromise performance [1, 7]. Moreover, simply reducing the batch size does not lower the FL communication costs, as clients still need to exchange the full model with the server. On the other hand, using a large batch size would exclude low-memory devices from participating in the FL process.
@@
Figure 1: Four key steps in a single FL communication round. (i) The server distributes a base global model $M$ to all participating clients. (ii) Each client $n\in[1,N]$ trains the model on its local dataset $D^{n}$ to produce a local model $M^{n}$. (iii) The local models are then transmitted back to the server. (iv) Finally, the server aggregates the received local models using weighted averaging to update the global model: $M\leftarrow\sum^{N}_{n=1}w^{n}M^{n}$ [5].
@@ 2.1 Federated Learning
A typical FL process involves a central coordinating server and a set of client devices, each storing private data [5]. The goal of FL is to train a global model $M$, by learning from the data residing on clients while ensuring their privacy. The FL objective can be expressed as:
@@
where $N$ is the number of clients, and $\mathcal{L}$ represents the local loss function. Here, $w^{n}=|D^{n}|/|D|$ is the weight assigned to the $n$-th client, with $|D^{n}|$ denoting the number of samples in the local dataset $D^{n}$, and $|D|$ denoting the total number of samples across all clients, i.e., ${D=\bigcup_{n=1}^{N}D^{n}}$. As shown in Fig. 1, a single FL communication round consists of four main steps, which are repeated for a total of $R$ rounds.
@@ 2.2 Self-supervised Learning
Self-supervised learning (SSL) offers a way to utilize unlabeled data for model training, which is often more readily available than labeled data. SSL generates its own supervisory signals to learn valuable representations from the unlabeled data. In this work, we primarily use MoCoV3 [11] as the SSL approach.111 Unlike in [11], we train the patch projection layer. As shown in Fig. 2, MoCoV3 features a Siamese network structure, consisting of an actively trained online branch, and a target branch. The online branch includes an encoder $F$, a projection head $H$, and a prediction head $P$. Meanwhile, the target branch is a moving-averaged copy of the online branch, consisting of a momentum encoder $F_{k}$ and a momentum projection head $H_{k}$.
Given an input sample $x$, it undergoes augmentation, creating two views, $x_{q}$ and $x_{k^{+}}$. These views are then fed into the online and target branches, respectively, to obtain representations $q$ and its corresponding positive pair, $k^{+}$. Likewise, negative pairs, $k^{-}$, are derived from other samples within the same batch through the target branch. The online branch is trained using the InfoNCE loss [12], defined as:
@@ 2.3 Federated Self-supervised Learning
While most studies on FL focus on supervised learning, this requires device owners to label their data—a labor-intensive and often impractical task. As a result, client devices often lack labeled data due to limitations in data collection, labeling resource constraints, and insufficient annotation expertise. Federated self-supervised learning (FedSSL) extends the applicability of FL to new domains where labeled data is scarce or difficult to acquire [13, 14, 15]. Many FedSSL studies build upon state-of-the-art SSL frameworks, such as SimCLR [1], MoCo [2], and BYOL [16]. These studies adapt SSL techniques to address specific FL tasks [17, 18] or tackle FL challenges at hand [19, 20, 21, 6]. In the context of FedSSL, the local dataset $D^{n}$ within a client $n$ is unlabeled, and thus, the local training process relies on a form of SSL loss to train the local model $M^{n}$. When using MoCoV3 as the SSL framework, the local model consists of an encoder $F^{n}$ and its accompanying MLP heads, $H^{n}$ and $P^{n}$.
@@ 2.4 Layer-wise Training
Layer-wise training was initially introduced in the context of deep belief networks and restricted Boltzmann machines [22, 23]. The original motivation behind its development was to address challenges associated with training deep neural networks, such as vanishing gradients. While its usage has become less common over the years, we believe it holds great potential to be rediscovered as a highly resource-efficient strategy within FL systems. Incrementally training a model one layer at a time offers a significant resource-saving advantage over end-to-end training. Despite its potential benefits, layer-wise training remains largely unexplored in FL, with very few related studies [24, 25].
The study in [26] investigates layer-wise SSL in the speech domain. Although it was intended for an on-device training scenario, the authors mainly examined it in a centralized setting. A short study [25] on federated layer-wise learning (FLL) introduces a depth dropout technique that randomly drops frozen layers during training to reduce resource overhead. However, FLL is only examined with a large number of communication rounds in each training stage (ranging from 4k to 12k), potentially placing significant strain on clients. Additionally, the exploration of FLL is limited in scope, missing comprehensive evaluations across diverse FL settings and benchmark datasets. In contrast to these studies, our work aims to offer a more thorough and in-depth exploration of layer-wise training through LW-FedSSL. One study [24] investigates the progressive training approach for FL known as ProgFed. Although similar to layer-wise training in gradually increasing the model depth, ProgFed [24] differs by training all existing layers at each stage. Since ProgFed primarily focuses on supervised learning tasks that require labeled data, the nature of progressive training within the FedSSL paradigm remains unexplored. To fill this gap, we introduce a progressive training strategy for FedSSL, referred to as Prog-FedSSL.
@@
13:         $\mathcal{L}\leftarrow\ell_{\text{con}}(q_{1},k_{2},\tau)+\ell_{\text{con}}(q_{2},k_{1},\tau)$
@@ 3.1 LW-FedSSL
Our proposed approach can be integrated with various SSL techniques, and we collectively refer to it as LW-FedSSL. Fig. 3 illustrates the local training process for the $n$-th FL client across different stages of LW-FedSSL. For an encoder $F$ with a total of $S$ layers, the training process can be generally divided into $S$ stages. Each stage $s\in[1,S]$ runs for a fixed number of FL communication rounds, $R_{s}$. The training process starts with an empty encoder $F_{0}$ (i.e., with no layers), and each stage $s$ sequentially adds a new layer $L_{s}$ to $F_{(s-1)}$, forming $F_{s}$ and increasing the encoder’s depth. During stage $s$, only the newly added layer $L_{s}$ is updated, while all prior layers [$L_{1}$, …, $L_{(s-1)}$] within $F_{s}$ are kept frozen, serving only for inference. Moreover, in each communication round, clients transmit only the active layer $L_{s}$ (along with the MLP heads, depending on the SSL technique used) to the server for aggregation. Likewise, the server only needs to broadcast the aggregated layer $L_{s}$ back to clients. The detailed procedure of LW-FedSSL using MoCoV3 as the SSL technique is shown in Algorithms 1 and 2. Algorithm 1 describes the server-side execution, which manages the incremental training stages $s\in[1,S]$ and the aggregation process. Each client executes Algorithm 2, performing local SSL training to update the active layer $L_{s}$.
@@
FedSSL end-to-end $F$ $[L_{1},\dots,L_{S}]$ $F$ $[L_{1},\dots,L_{S}]$ $[L_{1},\dots,L_{S}]$ $[L_{1},\dots,L_{S}]$ $[L_{1},\dots,L_{S}]$
LW-FedSSL layer-wise $S$ $F_{0}$ $F_{s}$ $[L_{1},\dots,L_{s}]$ $L_{s}$ $[L_{1},\dots,L_{(s-1)}]$ $L_{s}$ $L_{s}$ $L_{s}$
Prog-FedSSL progressive $S$ $F_{0}$ $F_{s}$ $[L_{1},\dots,L_{s}]$ $L_{s}$ $[L_{1},\dots,L_{s}]$ $[L_{1},\dots,L_{s}]$ $[L_{1},\dots,L_{s}]$
@@
3.2 Prog-FedSSL
Inspired by ProgFed [24], we also explore progressive training for federated self-supervised learning, denoted as Prog-FedSSL. The local training process of Prog-FedSSL at stage $s$ is shown in Fig. 4c, while a more comprehensive illustration across different stages is provided in Fig. 9 of A. Similar to layer-wise training, progressive training also starts with an empty encoder $F_{0}$ and sequentially adds a new layer $L_{s}$ to $F_{(s-1)}$ at the beginning of each stage $s$. However, in contrast to layer-wise training, which updates only a single layer $L_{s}$ at a time, progressive training updates all existing layers (i.e., [$L_{1}$, …, $L_{s}$]) during each stage $s$. In other words, no layers within $F_{s}$ are frozen at any stage. This distinction results in higher computational and communication costs for Prog-FedSSL compared to LW-FedSSL. Nonetheless, when compared with end-to-end training—where clients must train a full encoder with $S$ layers (i.e., [$L_{1}$, …, $L_{S}$]) at each round—Prog-FedSSL allows clients to train only $s$ layers $(s\leq S)$. Moreover, clients only need to exchange these $s$ layers with the server. In essence, Prog-FedSSL reduces both computational and communication costs by avoiding the training and exchanging of $(S-s)$ layers at each stage $s$ compared to end-to-end training. The detailed procedure of Prog-FedSSL using MoCoV3 as the SSL technique is presented in Algorithms 3 and 4 of A, with Algorithm 3 describing the server-side execution and Algorithm 4 describing the client-side execution.
@@ 4.1 Experimental setup
We use a ViT-Tiny [27] backbone with 12 transformer blocks as the encoder $F$. The projection head $H$ and the prediction head $P$ are implemented as a 3-layer MLP and a 2-layer MLP, respectively. For both $H$ and $P$, the hidden layer dimensions are set to 512, and the output dimension is set to 256.
@@
We use images from the COCO [28] dataset to construct the client datasets. The COCO dataset contains diverse natural scene images, which inherently introduce data heterogeneity. These images are uniformly distributed across 10 FL clients. To evaluate the performance, we use downstream datasets including CIFAR-10/100 [29], Tiny ImageNet [30], and Caltech-101 [31]. For all datasets, we use an image size of $32\times 32$ with a patch size of 4.
@@
BYOL [16]
BYOL [16] (Centralized) 87.69 66.26 44.84 62.32 65.28
@@
FedU [6] 8.94 (1.03×) 586 (1.00×) 8.40 (1.00×) 83.50 62.61 42.61 53.88 60.65
FedEMA [13] 8.96 (1.03×) 586 (1.00×) 8.40 (1.00×) 81.35 60.50 41.20 48.26 57.83
@@
SimCLR [1]
SimCLR [1] (Centralized) 87.49 65.64 43.54 61.12 64.45
@@
Training
[…]
Evaluation
After training, the encoder $F$ is retained, while the projection head $H$ and prediction head $P$ are discarded. A linear classifier head is added to the encoder $F$ and fine-tuned on the downstream datasets for 40 epochs, including a warmup period of 10 epochs. We set the batch size to 256 and use the AdamW optimizer with a base learning rate of 1e-3 and a weight decay of 1e-5. A cosine decay schedule is applied to the learning rate during training. We use the RandAugment function available in the PyTorch [34] framework for augmentation.

### v5

# Resource-efficient Layer-wise Federated Self-supervised Learning

@@ Abstract
Many studies integrate federated learning (FL) with self-supervised learning (SSL) to take advantage of raw data distributed across edge devices. However, edge devices often struggle with high computational and communication costs imposed by SSL and FL algorithms. With the deployment of more complex and large-scale models, these challenges are exacerbated. To tackle this, we propose Layer-Wise Federated Self-Supervised Learning (LW-FedSSL), which allows edge devices to incrementally train a small part of the model at a time. Specifically, in LW-FedSSL, training is decomposed into multiple stages, with each stage responsible for only a specific layer of the model. Since only a portion of the model is active for training at any given time, LW-FedSSL significantly reduces computational requirements. Additionally, only the active model portion needs to be exchanged between the FL server and clients, reducing communication overhead. This enables LW-FedSSL to jointly address both computational and communication challenges of FL client devices. It can achieve up to a $3.34\times$ reduction in memory usage, $4.20\times$ fewer computational operations (giga floating point operations, GFLOPs), and a $5.07\times$ lower communication cost while maintaining performance comparable to its end-to-end training counterpart. Furthermore, we explore a progressive training strategy called Progressive Federated Self-Supervised Learning (Prog-FedSSL), which offers a $1.84\times$ reduction in GFLOPs and a $1.67\times$ reduction in communication costs while maintaining the same memory requirements as end-to-end training. Although the resource efficiency of Prog-FedSSL is lower than that of LW-FedSSL, its performance improvements make it a viable candidate for FL environments with more lenient resource constraints.
@@ 1 Introduction
A significant portion of real-world data valuable for practical AI applications is distributed across edge devices, often subject to privacy constraints. Decentralized learning approaches like federated learning (FL) [konevcny2016federated, mcmahan2017communication] address these challenges by enabling privacy-preserving collaborative model training. FL systems operate by collecting trained model parameters from edge devices (a.k.a., clients) instead of their data, thereby preserving data privacy. This approach is particularly important for privacy-sensitive data such as medical records or financial information. However, much of the data on edge devices remains unlabeled, limiting the effectiveness of traditional supervised learning methods. To tackle this, self-supervised learning (SSL) strategies [chen2020simple, he2020momentum, chen2021exploring] allow models to learn from raw data by generating their own supervisory signals without requiring labels. Typically, SSL methods rely on centralized training schemes, which often involve exhaustive data collection to build a centrally stored dataset. Privacy concerns and the distributed nature of edge data make this centralized approach infeasible in many real-world scenarios. Consequently, federated self-supervised learning (FedSSL) has recently emerged as a promising approach that combines the strengths of FL and SSL, enabling models to leverage unlabeled data distributed across clients while preserving privacy [zhuang2021collaborative, li2023mocosfl, wang2023does].
Self-supervised learning (SSL), particularly with state-of-the-art models like Transformers [vaswani2017attention, dosovitskiy2021an], imposes substantial computational demands. Unfortunately, clients in an FL environment often operate with limited computational and communication resources. Many clients may lack the necessary resources to participate in a conventional end-to-end FedSSL process. As a result, data from these clients is excluded from the training process, potentially degrading overall model performance. Therefore, it is crucial to explore resource-efficient FedSSL approaches that enable every FL client to participate in collaborative training. A straightforward way to reduce computational demands on low-memory devices is to conduct the training with a smaller batch size. However, using a small batch size in SSL can diminish the quality of learned representations and compromise performance [chen2020simple, li2023mocosfl]. Moreover, simply reducing the batch size does not lower the FL communication costs, as clients still need to exchange the full model with the server. On the other hand, using a large batch size would exclude low-memory devices from participating in the FL process.
@@ 2.1 Federated Learning
A typical FL process involves a central coordinating server and a set of client devices, each storing private data [mcmahan2017communication]. The goal of FL is to train a global model $M$, by learning from the data residing on clients while ensuring their privacy. The FL objective can be expressed as:
@@
where $M^{*}$ is the optimal global model after training, $N$ is the number of clients, and $\mathcal{L}$ represents the local loss function. Here, $w^{n}=|D^{n}|/|D|$ is the weight assigned to the $n$-th client, with $|D^{n}|$ denoting the number of samples in the local dataset $D^{n}$, and $|D|$ denoting the total number of samples across all clients, i.e., $D=\bigcup_{n=1}^{N}D^{n}$}. As shown in Fig. 1, a single FL communication round consists of four main steps, which are repeated for a total of $R$ rounds.
Figure 1: Four key steps in a single FL communication round. (i) The server distributes a base global model $M$ to all participating clients. (ii) Each client $n\in[1,N]$ trains the model on its local dataset $D^{n}$ to produce a local model $M^{n}$. (iii) The local models are then transmitted back to the server. (iv) Finally, the server aggregates the received local models using weighted averaging to update the global model: $M\leftarrow\sum^{N}_{n=1}w^{n}M^{n}$ [mcmahan2017communication].
@@ 2.2 Self-supervised Learning
Self-supervised learning (SSL) offers a way to utilize unlabeled data for model training, which is often more readily available than labeled data. SSL generates its own supervisory signals to learn valuable representations from the unlabeled data. In this work, we primarily use MoCoV3 [9711302] as the SSL approach. [^1] Unlike in [9711302], we train the patch projection layer. As shown in Fig. 2, MoCoV3 features a Siamese network structure, consisting of an actively trained online branch, and a target branch. The online branch includes an encoder $F$, a projection head $H$, and a prediction head $P$. Meanwhile, the target branch is a moving-averaged copy of the online branch, consisting of a momentum encoder $F_{k}$ and a momentum projection head $H_{k}$.
Given an input sample $x$, it undergoes augmentation, creating two views, $x_{q}$ and $x_{k^{+}}$. These views are then fed into the online and target branches, respectively, to obtain representations $q$ and its corresponding positive pair, $k^{+}$. Likewise, negative pairs, $k^{-}$, are derived from other samples within the same batch through the target branch. The online branch is trained using the InfoNCE loss [van2018representation], defined as:
@@ 2.3 Federated Self-supervised Learning
While most studies on FL focus on supervised learning [mcmahan2017communication, li2020federated, Li_2021_CVPR], such approaches place the burden of maintaining labeled data on each device owner, which can be challenging in many real-world settings. Consequently, client devices often lack labeled data due to limitations in data collection, labeling resource constraints, and insufficient annotation expertise. Federated self-supervised learning (FedSSL) extends the applicability of FL to new domains where labeled data is scarce or difficult to acquire [zhuang2022divergence, saeed2020federated, ek2022federated]. Many FedSSL studies build upon state-of-the-art SSL frameworks, such as SimCLR [chen2020simple], MoCo [he2020momentum], and BYOL [grill2020bootstrap]. These studies adapt SSL techniques to address specific FL tasks [wu2022distributed, dong2021federated] or tackle FL challenges at hand [makhija2022federated, wu2021decentralized, zhang2023federated, zhuang2021collaborative]. In the context of FedSSL, the local dataset $D^{n}$ within a client $n$ is unlabeled, and thus, the local training process relies on a form of SSL loss to train the local model $M^{n}$. When using MoCoV3 as the SSL framework, the local model consists of an encoder $F^{n}$ and its accompanying MLP heads, $H^{n}$ and $P^{n}$.
@@ 2.4 Layer-wise Training
Layer-wise training was initially introduced in the context of deep belief networks and restricted Boltzmann machines [10.1162/neco.2006.18.7.1527, 10.5555/2976456.2976476]. The original motivation behind its development was to address challenges associated with training deep neural networks, such as vanishing gradients. While its usage has become less common over the years, we believe it holds great potential to be rediscovered as a highly resource-efficient strategy within FL systems. Incrementally training a model one layer at a time offers a significant resource-saving advantage over end-to-end training. Despite its potential benefits, layer-wise training remains largely unexplored in FL, with very few related studies [wang2022progfed, pengfei2023towards].
The study in [huo2021incremental] investigates layer-wise SSL in the speech domain. Although it was intended for an on-device training scenario, the authors mainly examined it in a centralized setting. A short study [pengfei2023towards] on federated layer-wise learning (FLL) introduces a depth dropout technique that randomly drops frozen layers during training to reduce resource overhead. However, FLL is only examined with a large number of communication rounds in each training stage (ranging from 4k to 12k), potentially placing significant strain on clients. Additionally, the exploration of FLL is limited in scope, missing comprehensive evaluations across diverse FL settings and benchmark datasets. In contrast to these studies, our work aims to offer a more thorough and in-depth exploration of layer-wise training through LW-FedSSL. One study [wang2022progfed] investigates the progressive training approach for FL known as ProgFed. Although similar to layer-wise training in gradually increasing the model depth, ProgFed [wang2022progfed] differs by training all existing layers at each stage. Since ProgFed primarily focuses on supervised learning tasks that require labeled data, the nature of progressive training within the FedSSL paradigm remains unexplored. To fill this gap, we introduce a progressive training strategy for FedSSL, referred to as Prog-FedSSL.
@@
13:   $\mathcal{L}\leftarrow\ell_{\text{con}}(q_{1},k_{2},\tau)+\ell_{\text{con}}(q_{2},k_{1},\tau)$ $\triangleright$ Local loss
@@ 3.1 LW-FedSSL
Our proposed approach can be integrated with various SSL techniques, and we collectively refer to it as LW-FedSSL. Fig. 3 illustrates the local training process for the $n$-th FL client across different stages of LW-FedSSL. For an encoder $F$ with a total of $S$ layers, the training process can be generally divided into $S$ stages. Each stage $s\in[1,S]$ runs for a fixed number of FL communication rounds, $R_{s}$. The training process starts with an empty encoder $F_{0}$ (i.e., with no layers), and each stage $s$ sequentially adds a new layer $L_{s}$ to $F_{(s-1)}$, forming $F_{s}$ and increasing the encoder’s depth. During stage $s$, only the newly added layer $L_{s}$ is updated, while all prior layers $L_{1}$, …, $L_{(s-1)}$] within $F_{s}$ are kept frozen, serving only for inference. Moreover, in each communication round, clients transmit only the active layer $L_{s}$ (along with the MLP heads, depending on the SSL technique used) to the server for aggregation. Likewise, the server only needs to broadcast the aggregated layer $L_{s}$ back to clients. The detailed procedure of LW-FedSSL using MoCoV3 as the SSL technique is shown in Algorithms 1 and 2. Algorithm 1 describes the server-side execution, which manages the incremental training stages $s\in[1,S]$ and the aggregation process. Each client executes Algorithm 2, performing local SSL training to update the active layer $L_{s}$.
@@
3.2 Prog-FedSSL
Inspired by ProgFed [wang2022progfed], we also explore progressive training for federated self-supervised learning, denoted as Prog-FedSSL. The local training process of Prog-FedSSL at stage $s$ is shown in Fig. 4c, while a more comprehensive illustration across different stages is provided in Fig. 11 of B. Similar to layer-wise training, progressive training also starts with an empty encoder $F_{0}$ and sequentially adds a new layer $L_{s}$ to $F_{(s-1)}$ at the beginning of each stage $s$. However, in contrast to layer-wise training, which updates only a single layer $L_{s}$ at a time, progressive training updates all existing layers (i.e., $L_{1}$, …, $L_{s}$]) during each stage $s$. In other words, no layers within $F_{s}$ are frozen at any stage. This distinction results in higher computational and communication costs for Prog-FedSSL compared to LW-FedSSL. Nonetheless, when compared with end-to-end training—where clients must train a full encoder with $S$ layers (i.e., $L_{1}$, …, $L_{S}$]) at each round—Prog-FedSSL allows clients to train only $s$ layers $(s\leq S)$. Moreover, clients only need to exchange these $s$ layers with the server. In essence, Prog-FedSSL reduces both computational and communication costs by avoiding the training and exchanging of (Ss)(S-s) layers at each stage $s$ compared to end-to-end training. The detailed procedure of Prog-FedSSL using MoCoV3 as the SSL technique is presented in Algorithms 3 and 4 of B, with Algorithm 3 describing the server-side execution and Algorithm 4 describing the client-side execution.
@@ 3.2 Prog-FedSSL
FedSSL end-to-end $F$ $L_{1},\dots,L_{S}]$ $F$ $L_{1},\dots,L_{S}] [$$L_{1},\dots,L_{S}] [$$L_{1},\dots,L_{S}] [$$L_{1},\dots,L_{S}]$
LW-FedSSL layer-wise $S$ $F_{0}$ $F_{s}$ $L_{1},\dots,L_{s}]$ $L_{s}$ $L_{1},\dots,L_{(s-1)}]$ $L_{s}$ $L_{s}$ $L_{s}$
Prog-FedSSL progressive $S$ $F_{0}$ $F_{s}$ $L_{1},\dots,L_{s}]$ $L_{s}$ $L_{1},\dots,L_{s}] [$$L_{1},\dots,L_{s}] [$$L_{1},\dots,L_{s}]$
@@ 4.1 Experimental setup
We use a ViT-Tiny [dosovitskiy2020image] backbone with 12 transformer blocks as the encoder $F$. The projection head $H$ and the prediction head $P$ are implemented as a 3-layer MLP and a 2-layer MLP, respectively. For both $H$ and $P$, the hidden layer dimensions are set to 512, and the output dimension is set to 256.
@@
We use images from the COCO [lin2014microsoft] dataset to construct the client datasets. The COCO dataset contains diverse natural scene images, which inherently introduce data heterogeneity. These images are uniformly distributed across 10 FL clients. To evaluate the performance, we use downstream datasets including CIFAR-10/100 [krizhevsky2009learning], Tiny ImageNet [le2015tiny], and Caltech-101 [li_andreeto_ranzato_perona_2022]. For all datasets, we use an image size of $32\times 32$ with a patch size of 4.
Training
The number of training stages is set to 12 (i.e., $S=12$). The total number of FL communication rounds $R$ is set to 180 for all approaches. For our proposed approaches, $R$ is evenly distributed across the stages $s\in[1,S]$}. This results in 15 rounds per stage (i.e., $R_{s}=R/S=15$}). Each FL client performs local training for 3 epochs using the AdamW optimizer, with a base learning rate of 1.5e-4 and a weight decay of 1e-5. The batch size is set to 512, and the learning rate is linearly scaled as $\text{base\_learning\_rate}\times\text{batch\_size}/\textit{256}$ [goyal2017accurate, krizhevsky2014one]. We set the momentum $\mu$ to 0.99 and the temperature $\tau$ to 0.05. Data augmentations for view creation include random resized crop, color jitter, grayscale, horizontal flip, Gaussian blur, and solarization.
Evaluation
After training, the encoder $F$ is retained, while the projection head $H$ and prediction head $P$ are discarded. A linear classifier head is added to the encoder $F$ and fine-tuned on the downstream datasets for 40 epochs, including a warmup period of 10 epochs. We set the batch size to 256 and use the AdamW optimizer with a base learning rate of 1e-3 and a weight decay of 1e-5. A cosine decay schedule is applied to the learning rate during training. We use the RandAugment function available in the PyTorch [Paszke_PyTorch_An_Imperative_2019] framework for augmentation.
@@
BYOL [grill2020bootstrap]
BYOL [grill2020bootstrap] (Centralized) 87.69 66.26 44.84 62.32 65.28
@@
FedU [zhuang2021collaborative] 8.94 (1.03×) 586 (1.00×) 8.40 (1.00×) 83.50 62.61 42.61 53.88 60.65
FedEMA [zhuang2022divergence] 8.96 (1.03×) 586 (1.00×) 8.40 (1.00×) 81.35 60.50 41.20 48.26 57.83
@@
SimCLR [chen2020simple]
SimCLR [chen2020simple] (Centralized) 87.49 65.64 43.54 61.12 64.45
@@
4.2 Performance and Resource Comparison
Table 2 presents a comparison of the resource requirements and performance across different training approaches. [^2] We report mean per-class accuracy for the Caltech-101 dataset. For resource requirements, we compare the computational and communication resources needed for a single FL client. Additionally, we demonstrate that our proposed methods, LW-FedSSL and Prog-FedSSL, can be seamlessly integrated with various SSL techniques. As shown in Table 2, we integrate MoCoV3 [9711302], BYOL [grill2020bootstrap], and SimCLR [chen2020simple] with our methods. For clarity, we group similar methods based on the SSL technique used. Fig. 5 plots the resource consumption for a client, illustrating computational resources (i.e., memory and FLOPs) across local epochs and communication cost (download + upload) across FL rounds. For FLOPs calculation, we consider only a single input sample.
The results in Table 2 show that LW-FedSSL significantly reduces resource requirements. Specifically, when using MoCoV3, LW-FedSSL achieves a $3.34\times$ reduction in memory usage, a $4.20\times$ fewer computational operations (GFLOPs), and a $5.07\times$ reduction in communication costs compared to conventional FedMoCoV3. Additionally, Table 2 demonstrates that LW-FedSSL also maintains comparable performance to its end-to-end training counterparts. While Prog-FedSSL is not as resource-efficient as LW-FedSSL, it still consumes fewer GFLOPs and communication resources compared to FedMoCoV3. Furthermore, Prog-FedSSL achieves better performance in most cases, suggesting that progressive training can also benefit self-supervised learning, aligning with the findings for supervised learning in [wang2022progfed]. Fig. 5 illustrates that the resource consumption for LW-FedSSL remains relatively flat across local epochs and FL rounds due to the focus on training a single layer at each stage. Meanwhile, the resource consumption for Prog-FedSSL gradually increases as the number of trainable layers grows with each stage, eventually matching that of conventional FedMoCoV3.
4.3 Worst-case Communication Scenario for LW-FedSSL
By default, we assume that clients only need to download the layer $L_{s}$ at each FL round for LW-FedSSL. This assumption holds only if all clients participate in every FL round without dropping out or no new clients join. Otherwise, new clients or those that previously dropped out would need to download the latest model $F_{s}$ (i.e., $L_{1},\dots,L_{s}]$) instead of just $L_{s}$. Therefore, we compare the worst-case communication cost for LW-FedSSL in Table 3, where a client downloads $F_{s}$ at every FL round. Note that client still needs to upload only $L_{s}$ after local training. The results show that even in the worst-case scenario, LW-FedSSL significantly reduces communication costs compared to end-to-end FedMoCoV3.
Comm. (GB)
FedMoCoV3 8.40 (1.00×)
LW-FedSSL (ours) 3.35 (0.40×)
Prog-FedSSL (ours) 5.04 (0.60×)